# Division of labor enables efficient collective decision-making under uncertainty

### v1

@@ 3.3 Uniform strategies degrade collective performance
A division of labor based on diverse risk tolerances is therefore crucial for balancing the costs and benefits of exploration (Fig. 3A). We next compare the performance of such optimal collectives to those with suboptimal strategy compositions. We quantify collective performance using two metrics: long-term average return $\rho$, normalized by the maximum possible return $\rho_{\max}$, and accuracy $\alpha$, defined as the time-averaged proportion of individuals committed to foraging when the environment is in the high-reward state (Fig. 3E).
When all thresholds are high, the population is overly cautious, since even relatively risk-tolerant individuals forage too infrequently to detect environmental improvements. As a result the collective responds slowly when good times return, reducing long-term average reward (Fig. 3C and E). In the extreme case where all individuals require excessive certainty to forage ($\theta_{n}>1/2$ for all $n$), unsuccessful foraging drives the collective belief below $1/2$, resulting in the cessation of foraging. In the absence of new observations, the belief relaxes toward complete uncertainty ($g=1/2$). Since all decision thresholds exceed this value, the collective remains inactive indefinitely.
Conversely, a collective with too many risk-takers fails in the opposite way: eager explorers can maintain high accuracy, but only by sustaining near-continuous sampling, so the marginal information gain of these scouts is outweighted by their high energetic cost, depressing long-term returns (Fig. 3D and E). In contrast, unstructured deviations from a shared risk-averse threshold primarily impair exploitation rather than information acquisition, desynchronizing the cautious majority and preventing coordinated surges of foraging when conditions are favorable (Fig. 3B and E).
Optimal collectives thus exhibit functional specialization, with a synchronized risk-averse majority sharing threshold $\theta_{c}$ responsible for coordinated exploitation, and a heterogeneous risk-tolerant minority sustaining exploration through a broad distribution of thresholds in large groups (Fig. 3 and SI Appendix, Fig. S4). Together, these results reveal a collective exploration–exploitation trade-off: heterogeneity among risk-tolerant individuals sustains exploration under uncertainty, while synchronization of the risk-averse majority maximizes returns once conditions are favorable. We next ask how ecological conditions reshape this optimal division of labor, when heterogeneity is favored, and when it collapses in favor of homogeneous strategies as the environment change from benign to volatile extremes.
3.4 Ecological extremes lead to a collapse in the division of labor
@@
Rapid environmental changes suppress heterogeneity, favoring homogeneous risk aversion over exploratory strategies (Fig. 4C). When environments fluctuate rapidly, maintaining risk-tolerant individuals becomes counterproductive as information gained through exploration quickly becomes obsolete. In this regime, the timescale of environmental changes is shorter than the timescale required for exploration and coordinated response. As a result, fast-changing environments favor conservative strategies that minimize exploration costs, leading to uniform populations of risk-averse individuals. By contrast, when environmental changes are slow and predictable relative to the timescale of exploration and response, information accumulates and remains reliable over long periods, so the group can profit from exploratory foraging in future decisions. Under these conditions, risk-tolerant individuals can repeatedly detect and confirm state changes, stabilizing collective belief and enabling timely collective commitment, favoring heterogeneous strategies.
@@ Discussion
We have shown that leaderless collectives can achieve optimal foraging performance through a division of labor based on heterogeneous risk tolerances. A small fraction of risk-tolerant individuals, whose number increases logarithmically with group size, maintains environmental awareness by exploring even when expected rewards are negative, while a synchronized risk-averse majority maximizes collective gains by committing only when conditions are favorable. This asymmetric structure reflects fundamental principles of collective information gathering: as individuals sample independently, the information value of additional observations diminishes due to redundancy, whereas exploration costs accumulate linearly with the number of scouts. As a result, only a small fraction of explorers is needed for accurate environmental tracking in large groups.
@@
Risk-tolerant individuals function as informative leaders in decentralized groups, guiding collective behavior by maintaining environmental awareness during periods of uncertainty [Strandburg-Peshkin2018, Couzin2005, King2009]. Whether such leadership is persistent or transient depends on its mechanistic origin. Stable phenotypic variation in energetic state or temperament can produce consistent leaders [Biro2008, King2009, Jolles2017], whereas probabilistic decision-making by otherwise homogeneous individuals allows leadership to rotate while preserving near-optimal collective performance [Couzin2011, Conradt2005, Jolles2020] (See SI Appendix, Section 4 for proof).
@@
While comparable divisions of labor can be imposed by centralized or dominance-based control in some systems [Conradt2005, Mech1999], we focused on how informational leadership can emerge in leaderless collectives via heterogeneity in exploration under uncertainty [Couzin2009, Hein2015]. A minority that samples when payoffs are ambiguous can distribute group sensing to regulate majority engagement, consistent with broader views of collective computation and information flow in animal groups [Sumpter2008]. This perspective suggests that if exploratory roles reflect stable phenotypic differences, individuals should show cross-context consistency in risk-taking and influence [Biro2008, Jolles2017, King2018]; if instead exploration arises from stochastic choice among otherwise similar individuals, leadership should rotate without loss of group-level performance [Couzin2011, Mann2018]. Correspondingly, altering exploration costs or environmental persistence is expected to shift the degree and stability of leadership and the strength of division of labor [Lima1990, Dall2005]. These predictions are experimentally accessible in social insects and other tractable systems where ecological timescales and costs can be manipulated and collective responses quantified [Seeley1994, Davidson2016, Dornhaus2006].
@@
A sublinearly increasing number of information leaders appears across diverse collective decision-making systems: informed minorities can effectively guide naive majorities in moving animal groups [Couzin2005]; a small fraction of hasty explorers can optimize group decision efficiency when observed by a large, deliberate majority [Karamched2020]; and a few scouts can reliably inform nest-site choices in social insects [Cronin2014]. Although leadership by risk-tolerant minorities has been studied experimentally in groups of fixed size [Reebs2000, Beekman2006, Dyer2008], how the number of leaders should scale with group size and how this scaling emerges from individual decision rules remains largely unexplored.
@@
The role of heterogeneity in group composition has been extensively studied using threshold-based models of division of labor, which typically assume symmetric, normally distributed response thresholds across individuals [Bonabeau1996, Bonabeau1998, Beshers2001, OSheaWheller2017, Ulrich2021]. However, the structure of threshold distributions itself has been largely overlooked. Skewness alone can generate complex relationships between group composition and collective performance. Division of labor in social insects can emerge from multiple interacting factors, beyond threshold variation [Ulrich2021], yet the contribution of the distribution structure remains unexplored. We hypothesize that natural groups may evolve skewed threshold compositions to optimize collective task performance under uncertainty.
@@
One limitation of our model is the assumption of perfect information sharing: all individuals instantaneously observe collective foraging outcomes. In many natural systems, however, not all potential foragers observe every returning individual, and information is transmitted locally or indirectly [Sumpter2005, Sumpter2008, Karamched2020]. Under such conditions, individuals may rely not only on explicit signals but also on implicit social information inferred from the behavior of others [Mann2018, Mann2020, Suganuma2025]. How partial information sharing and social inference reshape the optimal balance of division of labor remains an open question [Beshers2001]. We hypothesize that the core organizational structure identified here – a minority of explorers along with a synchronized majority – remains robust under imperfect communication, although achieving comparable performance may require a larger exploratory cohort to compensate for information loss.
@@
Our results raise unresolved questions about how finite energetic reserves constrain collective decision-making. While our model implicitly assumes unbounded accumulation of returns, biological systems operate under both lower and upper energetic limits. Near starvation, collectives may be forced into riskier strategies to avoid extinction. How such transitions are distributed across heterogeneous groups remains unclear. If risk heterogeneity reflects hunger-driven phenotypic variation, individuals with depleted reserves may become disproportionately risk-tolerant [Theraulaz1998], amplifying the influence of an informed minority during critical periods [King2009]. Conversely, when storage approaches saturation, continued foraging becomes inefficient, potentially suppressing exploration and diminishing the role of exploratory individuals altogether. Understanding how energetic constraints reshape the emergence, persistence, and influence of informed minorities is therefore essential for linking individual state, division of labor, and collective responsiveness of natural systems.
@@
In human society, the parallel to the risk-tolerant scout is thus not the state-sponsored adventurer like Zheng He or Magellan, but the self-motivated explorer: the garage inventor, the citizen scientist, the open-source builder, the high-risk entrepreneur, and the avant-garde artist. These individuals are driven by curiosity and passion and have lower thresholds for action than others. They often bear substantial personal costs, yet when they succeed, they spark movements that drive collective behavior at scale.
@@ 4 Methods
To understand how the optimal collective strategy, $\pi(g),$ changes with group size $N$, we perform asymptotic analysis of the Bellman equation in the limit of large $N$. This analysis characterizes the leading-order behavior of the optimal policy and shows that only a sublinear number of agents explore (forage) under unfavorable conditions, i.e. when $g<\theta_{c}$. In our model setup, this scaling is logarithmic in $N$.

### v2

@@ 3.3 Uniform strategies degrade collective performance
A division of labor based on diverse risk tolerances is therefore crucial for balancing the costs and benefits of exploration (Fig. 3A). We next compare the performance of such optimal collectives to those with suboptimal strategy compositions. We quantify collective performance using two metrics (Fig. 3E). The first is the long-term average return, $\rho$, normalized by the maximum possible return $\rho_{\max}=(\gamma_{+}-\lambda)/2$ achieved when all individuals forage only when conditions are favorable. The second is accuracy, $\alpha$, defined as the time-averaged proportion of individuals committed to foraging when the environment is in the high-reward state.
When all thresholds are high, the population is overly cautious, since even relatively risk-tolerant individuals forage too infrequently to detect environmental improvements. As a result, the collective responds slowly when good times return, reducing long-term average reward (Fig. 3C and E). In the extreme case where all individuals require excessive certainty to forage ($\theta_{n}>1/2$ for all $n$), unsuccessful foraging drives the collective belief below $1/2$, halting further exploration. In the absence of new observations, the belief relaxes back toward uncertainty ($g=1/2$). Since all decision thresholds exceed this value, the collective remains inactive indefinitely.
Conversely, a collective with too many risk-takers fails in the opposite way: eager explorers can maintain high accuracy, but only by sustaining near-continuous sampling. The marginal information gain of these explorers is outweighed by the energetic cost of foraging, depressing long-term returns (Fig. 3D and E). In contrast, unstructured deviations from a shared risk-averse threshold primarily impair exploitation rather than information acquisition, desynchronizing the cautious majority and preventing coordinated surges of foraging when conditions are favorable (Fig. 3B and E).
Optimal collectives thus exhibit functional specialization, with a synchronized risk-averse majority sharing threshold $\theta_{c}$ responsible for coordinated exploitation, and a heterogeneous risk-tolerant minority that sustains exploration through a broad distribution of thresholds in large groups (Fig. 3 and SI Appendix, Fig. S4). Together, these results reveal a collective exploration–exploitation trade-off: heterogeneity among risk-tolerant individuals sustains exploration under uncertainty, while synchronization of the risk-averse majority maximizes returns once conditions are favorable. We next ask how environmental conditions shape the optimal balance between explorers and exploiters, and identify when ecological pressures favor heterogeneous group compositions over uniform ones.
3.4 Ecological extremes suppress division of labor
@@
Rapid environmental changes suppress heterogeneity, favoring homogeneous risk aversion over exploratory strategies (Fig. 4C). When environments fluctuate rapidly, maintaining risk-tolerant individuals becomes counterproductive as information gained through exploration quickly becomes obsolete. In this regime, the timescale of environmental changes is shorter than the time needed to coordinate a response. As a result, fast-changing environments favor conservative strategies that minimize exploration costs, leading to uniform populations of risk-averse individuals. In contrast, when environmental changes are slow and predictable relative to the timescale of exploration and response, information accumulates and remains reliable over long periods, so the group can profit from exploratory foraging in future decisions. Under these conditions, risk-tolerant individuals can repeatedly detect state changes, stabilizing collective belief and enabling timely collective commitment, favoring heterogeneous strategies.
@@ Discussion
We have shown that leaderless collectives can achieve optimal foraging performance through a division of labor based on heterogeneous risk tolerances. A small fraction of risk-tolerant individuals, whose number increases logarithmically with group size, maintains environmental awareness by exploring even when expected rewards are negative, while a synchronized risk-averse majority maximizes collective gains by committing only when conditions are favorable. This asymmetric structure reflects fundamental principles of collective information gathering: as individuals sample independently, the information value of additional observations diminishes due to redundancy, whereas exploration costs accumulate linearly with the number of explorers, so only a small fraction is needed for accurate environmental tracking in large groups.
@@
Risk-tolerant individuals function as informative leaders in decentralized groups, guiding collective behavior by maintaining environmental awareness during periods of uncertainty [Strandburg-Peshkin2018, Couzin2005, King2009]. Whether such leadership is persistent or transient depends on its origin. Stable phenotypic variation in energetic state or temperament can produce consistent leaders [Biro2008, King2009, Jolles2017], whereas probabilistic decision-making by otherwise homogeneous individuals allows leadership to rotate while preserving near-optimal collective performance [Couzin2011, Conradt2005, Jolles2020] (See SI Appendix, Section 4 for proof).
@@
While comparable divisions of labor can be imposed by centralized or dominance-based control in some groups [Conradt2005, Mech1999], we focused on how informational leadership can emerge in leaderless collectives via heterogeneity in exploration preference under uncertainty [Couzin2009, Hein2015]. A minority that samples when payoffs are ambiguous can distribute group sensing to regulate majority engagement, consistent with broader views of collective computation and information flow in animal groups [Sumpter2008]. This perspective suggests that if exploratory roles reflect stable phenotypic differences, individuals should show cross-context consistency in risk-taking and influence [Biro2008, Jolles2017, King2018]; if instead exploration is due to stochastic choices among otherwise similar individuals, leadership should rotate without loss of group-level performance [Couzin2011, Mann2018]. Correspondingly, altering exploration costs or environmental persistence is expected to shift the degree and stability of leadership and the strength of division of labor [Lima1990, Dall2005]. These predictions are experimentally accessible in social insects and other tractable systems where ecological timescales and costs can be manipulated and collective responses quantified [Seeley1994, Davidson2016, Dornhaus2006].
@@
A sublinearly increasing number of information leaders appears in diverse collective decision-making systems: informed minorities can effectively guide naïve majorities in moving animal groups [Couzin2005], a small fraction of hasty explorers can optimize group decision efficiency when observed by a large, deliberate majority [Karamched2020], and a few scouts can reliably inform nest-site choices in social insects [Cronin2014]. Although leadership by risk-tolerant minorities has been studied experimentally in groups of fixed size [Reebs2000, Beekman2006, Dyer2008], how the number of leaders should scale with group size and how this scaling emerges from individual decision rules remains largely unexplored.
@@
The role of heterogeneity in group composition has been extensively studied using threshold-based models of division of labor, which typically assume symmetric, normally distributed response thresholds across individuals [Bonabeau1996, Bonabeau1998, Beshers2001, OSheaWheller2017, Ulrich2021]. However, the structure of threshold distributions itself has been largely overlooked. Skewness alone can generate complex relationships between group composition and collective performance. Division of labor in social insects can emerge from multiple interacting factors, beyond threshold variation [Ulrich2021], yet the contribution of the distribution structure remains unexplored. We hypothesize that natural groups may evolve skewed threshold compositions to optimize collective task performance.
@@
One limitation of our model is the assumption of perfect information sharing as all individuals instantaneously observe collective foraging outcomes. In many natural systems not all potential foragers observe every returning individual, and information is transmitted locally or indirectly [Sumpter2005, Sumpter2008, Karamched2020]. Such local transmission can generate modular organization, in which distinct subpopulations have unequal access to exploratory returns. Even when information is available, individuals may differ in perceptual or decision noise (e.g., noisy belief updates or variable thresholds) [Bogacz2006], resulting in heterogeneous beliefs and stochastic actions. We expect the qualitative division-of-labor strategy identified here to persist with such noise. Under such conditions, individuals may rely not only on explicit signals but also on implicit social information inferred from the behavior of others [Mann2018, Mann2020, Suganuma2025]. How partial information sharing and social inference reshape the optimal division of labor remains an open question [Beshers2001]. We hypothesize that the core organizational structure identified here – a minority of explorers along with a synchronized majority – remains robust under imperfect communication, although achieving comparable performance may require a larger exploratory cohort to compensate for information loss.
@@
Our results also raise questions about how finite energetic reserves constrain collective decision-making. While our model implicitly assumes unbounded accumulation of returns, biological systems operate under both lower and upper energetic limits. Near starvation, collectives may be forced into riskier strategies to avoid extinction. How such transitions are distributed across heterogeneous groups remains unclear. If risk heterogeneity reflects hunger-driven phenotypic variation, individuals with depleted reserves may become disproportionately risk-tolerant [Theraulaz1998], amplifying the influence of an informed minority during critical periods [King2009]. Conversely, when storage approaches saturation, continued foraging becomes unnecessary, potentially suppressing exploration and diminishing the role of exploratory individuals altogether. Understanding how energetic constraints reshape the emergence, persistence, and influence of informed minorities is therefore essential for linking individual state, division of labor, and collective responsiveness of natural systems.
@@
In human society, the parallel to the risk-tolerant explorer is thus not the state-sponsored adventurer like Zheng He or Magellan, but the self-motivated innovator: the garage inventor, the citizen scientist, the open-source builder, the high-risk entrepreneur, and the avant-garde artist. These individuals are driven by curiosity and passion and have lower thresholds for action than others. They often bear substantial personal costs, yet when they succeed, they spark movements that drive collective behavior at scale.
@@ 4 Methods
To understand how the optimal collective strategy, $\pi(g),$ changes with group size $N$, we perform asymptotic analysis of the Bellman equation in the limit of large $N$. This analysis characterizes the leading-order behavior of the optimal policy and shows that only a sublinear number of agents explore (forage) under unfavorable conditions, i.e., when $g<\theta_{c}$. In our model setup, this scaling is logarithmic in $N$.